OpenStudy (anonymous):

The vertices of a right triangle are (–6, 4), (0, 0), and (x, 4). What is the value of x?

OpenStudy (anonymous):

The vertices of a right triangle are (–6, 4), (0, 0), and (x, 4). What is the value of x?

OpenStudy (anonymous):

@StudyShark2000

OpenStudy (anonymous):

thank you for expaining most of my answers.

OpenStudy (anonymous):

This one is pretty tough. Frankly, I'm surprised you're having it assigned in middle school. If you sketch the points on paper, you'll see that there are two possibilities: -- 'x' could be ' 4 ', so the third point would be at (0, 4). Then the right triangle is upside down, with one of its legs up on top. Both legs are pieces of the 'x' and 'y' axes, and the hypotenuse is the line between the origin and (-6, 4). -- Another possibility: The third point could be up at y=4 and way over to the right of the y-axis, and the right angle could be at the origin (0, 0). Here's how to find it: If the third point is over there, and the right angle is at the origin, then the line from (-6, 4) to (0, 0) and the line from (x, 4) to (0, 0) are perpendicular. You know how to find the slopes of these two lines. - From (-6, 4) to (0, 0), the slope is (6 / -4) = - 3/2 . - From (x, 4) to (0, 0), the slope is (-x / -4) = x/4 . When two lines are perpendicular, their slopes are negative reciprocals ... (one slope) = -1 / (the other slope). So - 3/2 = -1 / (x/4) Multiply each side by -1 : 3/2 = 1 / (x/4) Multiply each side by x/4 : (3/2) (x/4) = 1 Multiply each side by 2/3 : x/4 = 2/3 Multiply each side by 4 : x = 8/3 = 2 and 2/3 . And no problem at all! Tag me on anymore if you need help with them.